OpenStudy (anonymous):

Michele_Laino A battery manufacturer randomly tests 500 batteries and finds that 3 are defective. How many defective batteries would be expected in a shipment of 12,000 batteries? A. 120 B. 72 C. 60 D. 36

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here the relative frequency of defective batteries, is: f=3/500 now, in order to get your answer you have to multiply that frequency by 12,000

OpenStudy (michele_laino):

what is: \[\large \frac{3}{{500}} \times 12000 = ...?\]

OpenStudy (anonymous):

72

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (anonymous):

yes :)

OpenStudy (michele_laino):

it is B